# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ 3.1. Description of the Algorithms
2     $m\leftarrow\lceil\frac{\sum_{i}\textsf{sz}_{i}}{\max_{i}\textsf{sz}_{i}}\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
3          With prob. $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{i}\textsf{sz}_{i}}$, sample $i\in\{1,\ldots,k\}$ if $S_{i}\neq\varnothing$ then
4               $\sigma\leftarrow S_{\ell}\!\boldsymbol{\cdot}\!\mathsf{dequeue}()$
5         else
6               break
7         if $\not\exists j<i,$ such that $O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma)$ then  $Y\leftarrow Y+1$
8    return $\frac{Y}{t}\cdot{\sum_{i}\textsf{sz}_{i}}$
@@ 2     $\eta^{\prime}\leftarrow\eta/4n$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ if $\ell=0$ then
6               ${P}^{\ell}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leftarrow\bigcup_{p\in{P}^{\ell}}\textsf{Pred}(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})$ // From the above set, remove all states that are not reachable from the initial state.
@@
The AppUnion subroutine. AppUnion ( Algorithm 2) takes as input $\epsilon,\delta$ and $\epsilon_{\textsf{sz}}$ and assumes access to sets $T_{1},\dots,T_{k}$. For each set $T_{i}$ the access is given in the form of membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements), and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$. Using these, Algorithm 2 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
The algorithm presented herein represents a modification of classical Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985). The proof methodology also shares similarities with  (Karp and Luby, 1985). For an arbitrary $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i$ such that $\sigma\in T_{i}$. And let set $\mathcal{U}_{1}$ be the set of all the $(\sigma,L(\sigma))$ pairs and let $\mathcal{U}_{2}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$. Consequently, the cardinality of the set $\mathcal{U}_{1}$ is $|\cup_{i=1}^{k}T_{i}|$, and that of $\mathcal{U}_{2}$ is $\sum_{i=1}^{k}|T_{i}|$. By drawing a sufficient number of samples from $\mathcal{U}_{2}$ and assessing the fraction belonging to $\mathcal{U}_{1}$, one can estimate $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$, which when multiplied by $|\mathcal{U}_{2}|$ provides a reliable estimate for the size of $\mathcal{U}_{1}$.
In the algorithm we take $t$ samples from $\mathcal{U}_{2}$. This involves initially sampling a set $T_{i}$ proportional to the size of the set (in line 2 ) and subsequently uniformly drawing a sample from $T_{i}$ (in line 2). As the exact sizes of the sets $T_{i}$ are unknown , estimates $\textsf{sz}_{i}$ are utilized for proportional sampling (in line 2), introducing an additional factor of error in the final estimate of $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$.This error can be mitigated by adjusting the value of $t$ accordingly.
Upon sampling an $i$ in Line 2, an element is chosen from the sample of the sets $S_{i}$. This mimics a random sample from $T_{i}$, given that $S_{i}$ already contains uniform samples from $T_{i}$. For each sampled index $i$ , a new element from $S_{i}$ is used and removed (Line 2). Initially, $S_{i}$ are sets are guaranteed to have a size of at least thresh, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop. If the algorithm ever needs more samples than $|\mathcal{U}_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence is demonstrated to be very low.
Finally, line 2 verifies whether the sample from $\mathcal{U}_{2}$ belongs to $\mathcal{U}_{1}$. It uses the oracles to the sets to check this. The variable $Y$ tallies the number of samples from $\mathcal{U}_{1}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{1}|/|\mathcal{U}_{2}|$. The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$, the estimate for $|\mathcal{U}_{2}|$.
Theorem 1 characterizes the formal guarantee of Algorithm 2.
Theorem 1.
Let $\epsilon,\delta>0$, and let $\Omega$ be some set. Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively. Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$. For each $i\leq k$, let $S_{i}$ be a sequence of samples (with possible repetitions) obtained by repeatedly sampling elements from $T_{i}$ uniformly with replacement such that $|S_{i}|\geq\texttt{thresh}$, where $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$ Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 2) outputs sz that satisfies:
${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\right]>1-\delta.$
The worst case number of calls to the oracles $O_{i}$ made by the algorithm is $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ and the rest of the computation takes a similar amount of time.
[…]
@@
  2. (2)
    For a state ${q}^{\ell}$, we will use $\mathsf{U}({q}^{\ell})$ to denote the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$. Our algorithms compute sequences $\mathsf{S}({q}^{\ell})$ of elements (strings) sampled from $\mathcal{L}({q}^{\ell})$. For arguing about correctness of our algorithms, we will in turn argue that the distribution of $\mathsf{S}({q}^{\ell})$ would be “similar" to that of $\mathsf{U}({q}^{\ell})$. Towards this, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. Formally, for any event $\mathcal{E},\mathcal{F}$ (in the context of Algorithm 3 and Algorithm 4) we will denote by ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp. ${\mathsf{Pr}}\!\left[\mathcal{E}\;|\;\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) the probability of the event $\mathcal{E}$ (resp. probability of the event $\mathcal{E}$, given event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$.
  3. (3)
  4. (4)
    For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, let us denote by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$. Here, $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. For ease of notation, we also define $\texttt{Accurate}\texttt{AppUnion}=\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$
  5. (5)
    Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 3). Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 3. Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
The following presents the correctness of Algorithm 3, while its proof is presented in Section 5.
Theorem 2.
Let $\gamma_{0},\beta$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy the following:
$\begin{array}[]{c}\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\quad\quad\beta\leq\frac{1}{4\cdot n^{2}},\quad\quad\text{ and,}\\ \forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\end{array}$
Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$, the following statements hold true
  1. (1)
    For each $w\in\mathcal{L}({q}^{\ell})$,
    ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
@@
    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ presented in unary

### 08 Apr 2024 v2

@@ 3.1. Approximating Union of Sets
2     $m\leftarrow\left\lceil\frac{\sum_{j}\textsf{sz}_{j}}{\max_{j}\textsf{sz}_{j}}\right\rceil$ $t\leftarrow\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot m}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ $Y\leftarrow 0$ repeat $t$ times:
3          Sample $i\in\{1,\ldots,k\}$ with probability $p_{i}=\frac{\textsf{sz}_{i}}{\sum_{j}\textsf{sz}_{j}}$ if $S_{i}\neq\varnothing$ then  $\sigma\leftarrow S_{i}\!\boldsymbol{\cdot}\!\mathsf{deque}(){}$ else  break if $\neg(\exists j<i,O_{j}\!\boldsymbol{\cdot}\!\mathsf{contains}(\sigma))$ then  $Y\leftarrow Y+1$
4    return $\frac{Y}{t}\cdot({\sum_{i}\textsf{sz}_{i}})$
@@
After sampling an index $i$, in Line 1 an element is chosen (and removed) from the sampled list $S_{i}$. In the case when the set $S_{i}$ is constructed by uniformly selecting elements from $T_{i}$, this step mimics drawing a random sample from the actual set $T_{i}$. Initially, each set (more precisely, list) $S_{i}$ is guaranteed to have a size of at least $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$, surpassing the expected number of samples required from $T_{i}$ during the $t$ iterations of the loop. If the algorithm ever needs more samples than $|S_{i}|$ during the $t$ iterations, it counts as an error, but the probability of this occurrence will be demonstrated to be very low. Finally, Line 1 verifies whether the sample from $\mathcal{U}_{\textsf{multiple}}$ belongs to $\mathcal{U}_{\textsf{unique}}$ by asking a membership question to the oracle $O_{i}$. If all the sampled sets $S_{i}$’s are constructed uniformly, then the variable $Y$ tallies the number of samples from $\mathcal{U}_{\textsf{unique}}$, and after $t$ iterations, $Y/t$ provides a estimation for $|\mathcal{U}_{\textsf{unique}}|/|\mathcal{U}_{\textsf{multiple}}|$. The final output is the product of this value with $(\sum_{i}\textsf{sz}_{i})$ (which is the estimate for $|\mathcal{U}_{\textsf{multiple}}|$).
We next present (in Theorem 1) the formal correctness statement of this algorithm using distribution entanglement (see Section 2.1). For this, we set the source random variable $X$ to be the one that corresponds to the product distribution of the input sampled sets $S_{1},\ldots,S_{k}$, while the target random variable $Y$ corresponds to the product distribution of ${\mathsf{U}}_{1},\ldots,{\mathsf{U}}_{k}$, where, ${\mathsf{U}}_{i}$, is the random variable that obeys the distribution obtained when constructing a subset of $T_{i}$ (of at least thresh many elements) by uniformly picking each element of $T_{i}$. Since the random variables $S_{1},\ldots,S_{k}$ are clear from context, we will use ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{1,\ldots,k}}\right]$ to denote the probability of event $\mathcal{E}$ in the resulting entanglement distribution. For well definedness, we require that ${\mathsf{Pr}}\!\left[S_{i}=\omega\right]>0$ for every subset $\omega\subseteq T_{i}$ (with $\omega\geq\texttt{thresh}$), and will ensure this each time we invoke this statement.
Theorem 1 ().
Let $\epsilon,\delta>0$, and let $\Omega$ be some set. Let $T_{1},T_{2},\ldots,T_{k}\subseteq\Omega$ be sets with membership oracles $O_{1},\ldots,O_{k}$ respectively. Let $\epsilon_{\textsf{sz}}\geq 0$ and let $\textsf{sz}_{1},\ldots,\textsf{sz}_{k}\in\mathbb{N}$ be such that for every $i\leq k$, we have $\frac{\left|T_{i}\right|}{(1+\epsilon_{\textsf{sz}})}\leq\textsf{sz}_{i}\leq(1+\epsilon_{\textsf{sz}})\left|T_{i}\right|$. Let $\texttt{thresh}=24\cdot\frac{(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\cdot\log(\frac{4k}{\delta})$. For each $i\leq k$, let $S_{i}$ be some sequence of samples (with possible repetitions), obtained according to some distribution such that $|S_{i}|\geq\texttt{thresh}$ and for every $\omega\subseteq T_{i}$ with $|\omega|\geq\texttt{thresh}$, ${\mathsf{Pr}}\!\left[S_{i}=\omega\right]>0$. Then, on input $\epsilon_{\textsf{sz}},(O_{1},S_{1},\textsf{sz}_{1}),\ldots,(O_{k},S_{k},\textsf{sz}_{k})$, the algorithm AppUnion with parameters $\epsilon,\delta$ (Algorithm 1) outputs sz that satisfies:
${\mathsf{Pr}}\!\left[\frac{|\cup_{i=1}^{k}T_{i}|}{(1+\epsilon)(1+\epsilon_{\textsf{sz}})}\leq\left|\textsf{sz}\right|<(1+\epsilon)(1+\epsilon_{\textsf{sz}})|\cup_{i=1}^{k}T_{i}|\,;{{\mathsf{U}}_{1,\ldots,k}}\right]>1-\delta.$
The algorithm makes $O\left(k\cdot(1+\epsilon_{\textsf{sz}})^{2}\cdot\frac{1}{\epsilon^{2}}\cdot\log\big{(}\frac{k}{\delta}\big{)}\right)$ number of membership calls to the oracles $O_{1},\ldots,O_{k}$ in the worst case and the rest of the computation takes a similar amount of time.
3.2. Sampling Subroutine
[…]
@@ 2     $\eta^{\prime}\leftarrow\eta/4n$ $\beta^{\prime}\leftarrow(1+\beta)^{\ell-1}-1$ if $\ell=0$ then
  • For a state ${q}^{\ell}$, we use $\mathsf{U}({q}^{\ell})$ to represent the random variable denoting the sequence of samples (with possible repetitions) constructed by repeatedly sampling $|\mathsf{S}({q}^{\ell})|$ number of elements uniformly from $\mathcal{L}({q}^{\ell})$. Let ${\sf S}_{\leq i}$ be the random variable corresponding to the product of $\{\mathsf{S}({q}^{j})\}_{q\in Q,0\leq j\leq i}$. Also, let ${\mathsf{U}}_{\leq i}$ represent the random variable obtained by taking the ordered profuct of all $\{\mathsf{U}({q}^{j})\}_{q\in Q,0\leq j\leq i}$. In our algorithms, we will use ${\mathsf{Pr}}\!\left[\mathcal{E};{{\mathsf{U}}_{\leq i}}\right]$ (resp. ${\mathsf{Pr}}\!\left[\mathcal{E}|\mathcal{F};{{\mathsf{U}}_{\leq i}}\right]$) to denote the probability (resp. conditional probability) of $\mathcal{E}$ (resp. $\mathcal{E}$ conditioned on $\mathcal{F}$) when ${\sf S}_{\leq i}$ is distribution entangled to ${\mathsf{U}}_{\leq i}$.
  • For each $1\leq i\leq\ell$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$, by $\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ we denote the event that the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}^{\text{th}}$ call to AppUnion corresponding to level $i$ returns a value $\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ that satisfies 1($\frac{1}{(1+\beta)^{i}}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|\leq\textsf{sz}^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\leq(1+\beta)^{i}\cdot|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}})|$. Here, $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. For ease of notation, we also define $\texttt{Accurate}\texttt{AppUnion}=\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$
  • Let $\mathtt{Fail}_{1}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned because $\varphi>1$ at the time of return (Line 2). Let $\mathtt{Fail}_{2}$ be the event that $\mathtt{ret}_{\textnormal{{sample}}}=\bot$ is returned at Line 2. Finally, let $\mathtt{Fail}=\mathtt{Fail}_{1}\uplus\mathtt{Fail}_{2}$.
The following presents the correctness of Algorithm 2, while its proof is presented in Section 4.2.
Theorem 2 ().
Let $\gamma_{0}=\frac{2}{3e}\cdot\frac{1}{\mathsf{N}({q}^{\ell})},\beta\leq\frac{1}{4\cdot n^{2}}$ and $\{\mathsf{S}({p}^{i})\}_{1\leq i<\ell,p\in Q}$ satisfy
 $\forall\;1\leq i\leq\ell,\forall p\in Q,\;\;|\mathsf{S}({p}^{i})|\geq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right).$
Then for the trial corresponding to the invocation $\textnormal{{sample}}(\ell,\{{q}^{\ell}\},\lambda,\gamma_{0},\beta,\eta)$ the following hold
@@
    For each $w\in\mathcal{L}({q}^{\ell})$, ${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\;\;\big{|}\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=\gamma_{0}$
@@
    ${\mathsf{Pr}}\!\left[\texttt{Accurate}\texttt{AppUnion}\mid\texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-\frac{\eta}{2}$
3.3. Main Algorithm
Input : NFA $\mathcal{A}$ with $m$ states, $n\in\mathbb{N}$ in unary